OpenStudy (anonymous):

Solve the following system of equations. Please show your work to receive full credit. x - y = 16 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice

OpenStudy (ivettef365):

you can add as follows: x - y = 16 2x +y = 2 ==========

OpenStudy (ivettef365):

do you now how to add the equation?

OpenStudy (anonymous):

Unfortunally I do not.

OpenStudy (ivettef365):

ok this is how x -y = 16 2x +y = 2 ========== 3x = 18

OpenStudy (ivettef365):

now divide by 3 and you have 3x 18 --- = ---- 3 3

OpenStudy (ivettef365):

so what is x equal to ?

OpenStudy (anonymous):

x = 6

OpenStudy (ivettef365):

good, now to find the value of y just substitute the 6 for x on the equation as follows: x - y = 16 6 - y = 16 can you solve for y

OpenStudy (anonymous):

I don't think so. The way I look at it it be negative if i get a total of it :/

OpenStudy (anonymous):

Oh wait it be -10 then?

OpenStudy (ivettef365):

correct, so now you solved the system of equation (6,-10)

OpenStudy (anonymous):

Question

OpenStudy (ivettef365):

you solved it by addition

OpenStudy (anonymous):

x -y = 16 2x +y = 2 ========== 3x = 18 <----- for this one how did u get 3x

OpenStudy (anonymous):

I never understood that one

OpenStudy (ivettef365):

ok x + 2x = 3x -y +y = 0 (left it blank) 16 + 2 = 18

OpenStudy (anonymous):

so x would be 1?

OpenStudy (ivettef365):

right x = 1x

OpenStudy (anonymous):

now that make sense. ty

OpenStudy (ivettef365):

yw :)

OpenStudy (anonymous):

umm for step 2

OpenStudy (anonymous):

How can i explain it ?

OpenStudy (ivettef365):

you used the addition method

OpenStudy (anonymous):

Oh i see

OpenStudy (anonymous):

Addition method seems alot easy to do didnt feel it needed so much work to slove it.

OpenStudy (ivettef365):

since you were able to cancel out the y right away

OpenStudy (anonymous):

Well thank you so much ^^ I got it now

OpenStudy (ivettef365):

no problem, anytime